# Gravitational dressing: from the crossed product to more general algebraic and mathematical structure

### v2

@@
An important question is extending such relational constructions to the quantum context. [^3] In the quantum context, the general relational constructions may also be described in the terminology of quantum reference frames[23, 24, 25]. This paper will focus on the case with $X^{A}[g_{\mu\nu}]$, independent of other fields, which we refer to as gravitationally dressed, in contrast[12, 13] to $X^{A}[\phi^{I}]$, the field dressed case, or a more general mixed observables. The gravitationally dressed case arguably represents the minimal structure needed in gravity[13], since it involves only gravitational degrees of freedom; we expect the corresponding quantum operators to create/annihilate quantum field configurations together with their gravitational fields.
@@
$H^{0}_{ij}(x^{\prime},x)=H^{0}_{S,ij}(x^{\prime})+\Delta H^{0}_{ij}(x^{\prime},x)\ ,$ (28)
where $\kappa H^{0}_{S,ij}(x^{\prime})$ is, for $y>0$, the metric deformation corresponding to a unit-mass deformation of Schwarzschild, $\partial_{M}q_{ij}$ from (26), and vanishes for $y<0$. This can be thought of as providing a standard dressing associated with the horizon; specifically, when acted on with the derivative operators of (10), it will have a source localized at the horizon, $y=0$. The remaining term $\Delta H$ can be thought of as providing a “dressing to the horizon,” together with an analog of the angular momentum piece in (17). For the full dressing one also requires $G^{ijk}$; $H^{k}_{ij}$ and $G^{ij0}$ can be taken to vanish due to vanishing of the background extrinsic curvature $K_{ij}$ of the extremal slices.
@@
$V^{0}(x)$ $=$ $V^{0}_{S}+\Delta V^{0}(x)=\frac{\kappa}{2}\int d^{3}x^{\prime}H^{0}_{S,ij}(x^{\prime})p^{ij}(x^{\prime})+\frac{\kappa}{2}\int d^{3}x^{\prime}\Delta H^{\mu}_{ij}(x^{\prime},x)p^{ij}(x^{\prime})$ (29)
@@
$\int d^{3}x\sqrt{q}n^{\nu}V^{\mu}(x)T_{\mu\nu}=V_{S}^{0}\int d^{3}x\sqrt{q}\xi^{\mu}n^{\nu}T_{\mu\nu}+\int d^{3}x\sqrt{q}n^{\mu}\left[\Delta V^{0}T_{0\mu}+V^{k}(x)T_{k\mu}\right]\ .$ (31)
This clearly illustrates a reduction to the crossed product as a truncation of the more complete gravitational dressing: if the last terms of eq. (31) are dropped, what remains is $V_{S}^{0}P_{t}$. As noted, $P_{t}$ is an automorphism – specifically the modular automorphism associated to the Hartle-Hawking vacuum – of the subalgebra ${\cal A}_{L}$ of operators inside the horizon, and since the spacetime is asymptotically flat, the standard dressing $V^{0}_{S}$ is again conjugate to the corresponding ADM momentum,
$P_{0}^{ADM},V^{0}_{S}]=i\ .$ (32)
@@
Various observations can now be made. First is the question of what algebra or states enter the crossed product construction. In the way it has been described here, the states and algebra correspond to excitations to the left of the horizon of the eternal black hole, dressed to the right. This algebra was discussed in [17], who also discuss the related algebras, which also can be constructed via the standard dressing, namely ${\cal A}_{R}$, “dual” to right operators dressed to the right, or in an analogous construction with a left dressing, ${\cal A}_{L}$ “dual” to left operators dressed to the left. Switching between right and left truncated dressing can be accomplished by subtracting $\partial_{M}q_{ij}$ from $\kappa H^{0}_{ij}$, as in (28). In an alternate construction, one could begin with right operators fully dressed to the right, and then make the same subtraction, so that the dressing has nonzero commutator with the left rather than the right ADM hamiltonian; while perhaps more contrived, this also appears to yield a crossed product without truncation. [^7] I thank Xi Dong for a discussion on this. Note also that $V^{0}_{S}$ provides an actual gravitational realization of the “time-shift” operator described in [17]; indeed, compare (3) and (6).

### v3

@@
An important question is extending such relational constructions to the quantum context. [^3] In the quantum context, the general relational constructions may also be described in the terminology of quantum reference frames[23, 24, 25]. Related constructions in the language of the “dressing field method” are described in ref. [26], and references therein. This paper will focus on the case with $X^{A}[g_{\mu\nu}]$, independent of other fields, which we refer to as gravitationally dressed, in contrast[12, 13] to $X^{A}[\phi^{I}]$, the field dressed case, or a more general mixed observables. The gravitationally dressed case arguably represents the minimal structure needed in gravity[13], since it involves only gravitational degrees of freedom; we expect the corresponding quantum operators to create/annihilate quantum field configurations together with their gravitational fields.
@@
$H^{0}_{ij}(x^{\prime},x)=H^{0}_{RS,ij}(x^{\prime})+\Delta H^{0}_{ij}(x^{\prime},x)\ ,$ (28)
where $\kappa H^{0}_{RS,ij}(x^{\prime})$ is, for $y>0$, the metric deformation corresponding to a unit-mass deformation of Schwarzschild, $\partial_{M}q_{ij}$ from (26), and vanishes for $y<0$. This can be thought of as providing a (right) standard dressing associated with the horizon; specifically, when acted on with the derivative operators of (10), it will have a source localized at the horizon, $y=0$. The remaining term $\Delta H$ can be thought of as providing a “dressing to the horizon,” together with an analog of the angular momentum piece in (17). For the full dressing one also requires $G^{ijk}$; $H^{k}_{ij}$ and $G^{ij0}$ can be taken to vanish due to vanishing of the background extrinsic curvature $K_{ij}$ of the extremal slices.
@@
$V^{0}(x)$ $=$ $V^{0}_{RS}+\Delta V^{0}(x)=\frac{\kappa}{2}\int d^{3}x^{\prime}H^{0}_{S,ij}(x^{\prime})p^{ij}(x^{\prime})+\frac{\kappa}{2}\int d^{3}x^{\prime}\Delta H^{\mu}_{ij}(x^{\prime},x)p^{ij}(x^{\prime})$ (29)
@@
$\int d^{3}x\sqrt{q}n^{\nu}V^{\mu}(x)T_{\mu\nu}=V_{RS}^{0}\int d^{3}x\sqrt{q}\xi^{\mu}n^{\nu}T_{\mu\nu}+\int d^{3}x\sqrt{q}n^{\mu}\left[\Delta V^{0}T_{0\mu}+V^{k}(x)T_{k\mu}\right]\ .$ (31)
This clearly illustrates a reduction to the crossed product as a truncation of the more complete gravitational dressing: if the last terms of eq. (31) are dropped, what remains is $V_{RS}^{0}P_{t}$. As noted, $P_{t}$ is an automorphism – specifically the modular automorphism associated to the Hartle-Hawking vacuum – of the subalgebra ${\cal A}_{L}$ of operators inside the horizon, and since the spacetime is asymptotically flat, the standard dressing $V^{0}_{S}$ is again conjugate to the corresponding (right) ADM momentum,
$P_{R0}^{ADM},V^{0}_{RS}]=i\ .$ (32)
@@
Various observations can now be made. First is the question of what algebra or states enter the crossed product construction. In the way it has been described here, the states and algebra correspond to excitations to the left of the horizon of the eternal black hole, dressed to the right. This and related algebras were discussed in [17]; the related algebras, which also can be constructed via the standard dressing, include ${\cal A}_{R}$, “dual” to right operators dressed to the right, or in an analogous construction with a left dressing, ${\cal A}_{L}$ “dual” to left operators dressed to the left. Switching between right and left truncated dressing can be accomplished by subtracting $\partial_{M}q_{ij}$ from $\kappa H^{0}_{ij}$, as in (28). In an alternate construction, one could begin with right operators fully dressed to the right, and then make the same subtraction, so that the dressing has nonzero commutator with the left rather than the right ADM hamiltonian; while perhaps more contrived, this also appears to yield a crossed product without truncation. [^7] I thank Xi Dong for a discussion on this. Note also that $V^{0}_{RS}$ provides a realization in gravitational varaibles of the “time-shift” operator described in [17]; indeed, compare (3) and (6).